OpenStudy (anonymous):

The lateral area of a cone is 480 pi cm2. The radius is 40 cm. Find the slant height.

OpenStudy (kropot72):

The formula for the lateral area of a cone is \[lateral\ area=\pi \times r \times l\] where l = slant height The formula can be rearranged to find the slant height in terms of the lateral area and the radius as follows: \[l=\frac{lateral\ area}{\pi \times r}\] Now you just need to plug the given values of lateral area and radius into the latter equation to find the slant height.

OpenStudy (kropot72):

\[l=\frac{480\times \pi}{\pi \times 40}cm=you\ can\ calculate\]

OpenStudy (anonymous):

Um... I got 19, 200

OpenStudy (kropot72):

\[l=\frac{480\times \pi}{\pi \times 40}=\frac{480}{40}cm\] What do you get when 480 is divided by 40?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

is that it or is there something next???

OpenStudy (kropot72):

The slant height is 12 cm.